A company has $89 million in outstanding bonds, and 10 million shares of stock currently trading at $34 per share.The bonds pay an annual coupon rate of 8% and is trading at par. The company's beta is 1.2, its tax rate is 40%, the risk-free rate is 2%, and the market risk premium is 5%. What is this firm's WACC?
Answer #1

Solution:

Calculation of cost of equity :

Cost of equity as per Capital Asset Pricing Model is calculated using the following formula :

RE = RF + β ( RM - RF )

Where

RE = Cost of equity    ; RF = Risk free rate of return   ; β = Beta of the stock ; ( RM – RF ) = Market risk Premium

As per the information given in the question we have

RF = 2 %   ; ( RM - RF ) = Market Risk Premium = 5 %   ; β = 1.2

Applying the above values in the formula we have

= 2 % + ( 1.2 * 5 % )

= 2 % + 6 % = 8 %

Thus the cost of equity = 8 %        

As per the information given in the question            

Market value of the Equity shares or Common stock = No. of shares outstanding * Market price per share

= $ 10 Million shares * $ 34 = $ 340 Million

Face value of the bonds = $ 89 million

Thus Weight of equity shares or Common stock = [ 340 / ( 340 + 89 ) ] = 340 / 429 = 0.792541

Thus Weight of corporate bonds = [ 89 / ( 340 + 89 ) ] = 89 / 429 = 0.207459

Cost of equity = 8 %   ; Cost of corporate bonds = cost of debt = 8 %

Calculation of WACC :

The formula for calculating the weighted average cost of capital is =

WACC = [ Ke * We ] + [ ( Kd * ( 1- t ) ) * Wd ]

Ke = Cost of equity ; We = Weight of equity ; Kd = Cost of debt    ; t = Income tax rate ; Wd = Weight of debt

As per the information available in the question we have

Ke = 8 % = 0.08   ; We = 0.792541 ;   Kd = 8 % = 0.08 ; t = 40 % = 0.40   ; Wd = 0.207459

Applying the above values in the formula we have

= [ 0.08 * 0.792541 ] + [ (0.08 * ( 1 – 0.40 ) ) * 0.207459 ]

= [ 0.08 * 0.792541 ] + [ (0.08 * 0.60 * 0.207459 ]

= [ 0.063403 + 0.009958   ]

= 0.073361

= 0.0734 ( when rounded off to four decimal places)

= 7.34 %

Thus the WACC of the firm is = 7.34 %